### v2

@@ 6 Numerical Results
We implemented Algorithm 1 and Algorithm 2 and used them on several classes of combinatorial structures shown in Table 1 and checked the results against the Online Encyclopedia of Integer Sequences (OEIS). [^11] https://oeis.org/ In the evaluation, we counted the structures described below, where we use
@@ 4 $16uv^{6}+15uv^{8}+6uv^{10}+uv^{12}+15u^{2}v^{4}+4u^{2}v^{6}+6u^{3}v^{2}+u^{4}$ 16 38
Both OEIS sequence A000272 and $\mathsf{WFOMC}(\Psi_{ug}\land tree(R),n,\mathds{1},\mathds{1})$ count the number of trees on $n$ labeled nodes. By Theorem 3.10, we compute $tr_{n}=[uv^{2(n-1)}]f_{n}(u-1,v;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $tr_{n}$ matches the sequence A000272.
@@
Both OEIS sequence A001858 and $\mathsf{WFOMC}(\Psi_{ug}\land forest(R),n,\mathds{1},\mathds{1})$ count the number of forests on $n$ labeled nodes. By Theorem 3.11, we compute $fo_{n}=\sum_{i\in[n]}[u^{i}v^{2(n-i)}]f_{n}(u-1,v;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $fo_{n}$ matches the sequence A001858.
@@
By Theorem 4.10 and [MBS23, Proposition 14], we compute $dt_{n}=(-1)^{n}\cdot g_{n}(0,-2;\Psi_{dt},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $dt_{n}$ matches the sequence A000169.
@@ 7 Conclusion
A natural direction for future work involves identifying additional domain-liftable axioms with our proposed polynomials. Additionally, within the field of enumerative combinatorics, using the proposed polynomials to discover intriguing integer sequences, similar to those found in the OEIS, and in [SJT+23], also holds a substantial interest. Finally, from the perspective of complexity theory, our approach offers a unified framework for studying the upper bounds, i.e., the domain-liftability, of the model counting problem with axioms. However, the lower bounds for this problem remain open. It is worth exploring whether there exists a general criterion for determining the hardness of the model counting problem with axioms.
@@ Acknowledgements
Ondřej Kuželka’s work was supported by Czech Science Foundation project 24-11820S (“Automatic Combinatorialist”). Yuyi Wang’s work was partially supported by the Natural Science Foundation of Hunan Province, China (Grant No. 2024JJ5128).

### v3

@@ 6 Numerical Results
We implemented Algorithm 1 and Algorithm 2111111The source code can be found in https://github.com/l2l7l9p/Polynomials-for-WFOMC. and used them on several classes of combinatorial structures shown in Table 1 and checked the results against the Online Encyclopedia of Integer Sequences (OEIS). [^12] https://oeis.org/ In the evaluation, we counted the structures described below, where we use
@@ 4 $16uv^{6}+15uv^{8}+6uv^{10}+uv^{12}+15u^{2}v^{4}+4u^{2}v^{6}+6u^{3}v^{2}+u^{4}$ 16 38
Both OEIS sequence A000272 and $\mathsf{WFOMC}(\Psi_{ug}\land tree(R),n,\mathds{1},\mathds{1})$ count the number of trees on $n$ labeled nodes. By Theorem 3.10, we compute $tr_{n}=[uv^{2(n-1)}]\hat{f}_{n}(u-1,v;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $tr_{n}$ matches the sequence A000272.
@@
Both OEIS sequence A001858 and $\mathsf{WFOMC}(\Psi_{ug}\land forest(R),n,\mathds{1},\mathds{1})$ count the number of forests on $n$ labeled nodes. By Theorem 3.11, we compute $fo_{n}=\sum_{i\in[n]}[u^{i}v^{2(n-i)}]\hat{f}_{n}(u-1,v;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $fo_{n}$ matches the sequence A001858.
@@
By Theorem 4.10 and [WFOMC-axioms, Proposition 14], we compute $dt_{n}=(-1)^{n}\cdot g_{n}(0,-2;\Psi_{dt},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $dt_{n}$ matches the sequence A000169.
@@ 7 Conclusion
A natural direction for future work involves identifying additional domain-liftable axioms with our proposed polynomials. Additionally, within the field of enumerative combinatorics, using the proposed polynomials to discover intriguing integer sequences, similar to those found in the OEIS, and in [fluffy], also holds a substantial interest. Finally, from the perspective of complexity theory, our approach offers a unified framework for studying the upper bounds, i.e., the domain-liftability, of the model counting problem with axioms. However, the lower bounds for this problem remain open. It is worth exploring whether there exists a general criterion for determining the hardness of the model counting problem with axioms.
@@ Acknowledgements
Ondřej Kuželka’s work was supported by Czech Science Foundation project 24-11820S (“Automatic Combinatorialist”). Yuanhong Wang’s work was supported by National Natural Science Foundation of China (No.62506141). Yuyi Wang’s work was partially supported by the Natural Science Foundation of Hunan Province, China (Grant No. 2024JJ5128).